OpenStudy (anonymous):

A package contains 9 candy canes, 6 of which are cracked. If 2 are selected, find the probability of getting no cracked candy canes. 0.3852 0.1938 0.2870 0.0833

OpenStudy (anonymous):

\[\frac{3}{9}\times \frac{2}{8}\]

OpenStudy (anonymous):

Take the ratio of the number of different ways you can pick 2 candy canes from the set of 3 uncracked ones, and the total number of ways you can pick 2 candy canes from a set of 9.

OpenStudy (anonymous):

means "didn't get one on the first selection AND didn't get one on the second"